# FedRTS: Federated Robust Pruning via Combinatorial Thompson Sampling

### v1

@@ 3.3 Theoretical Analysis
For simplifications, we omit the inter loop of FedRTS to isolate the TSAdj mechanism and consider only semi-outcomes, i.e., ignore the outcomes $X_{t,i}$ that $i\notin S_{t}$. We also consider the following assumption and approximation:
@@
(Independent Magnitude) Magnitude of model weights are mutually independent.
@@
(Threshold Approximation) The discriminative function $h(\cdot)$ in Eq. 4 and  5 can be approximated by a threshold $\sigma$, i.e., $h(i,x,\kappa)\approx\mathbf{1}_{x\geq\sigma}$.
Assumption 3.1 aligns with standard magnitude pruning practices, while Assumption 3.2 holds empirically, as weights often exhibit a Gaussian-like distribution under L2 regularization and normalization techniques. Therefore, under assumptions 3.1 and 3.2, the outcomes $X_{t,i}$ become:
@@
which are mutually independent. We also include the Lipschitz assumption 3.3 as in previous work (Kong et al., 2021).
Assumption 3.3.
(Lipschitz continuity) $\exists L\in\mathbb{R},\forall S,\mu,\mu^{\prime}$ satisfies $|r(S,\mu)-r(S,\mu^{\prime})|\leq L||\mu-\mu^{\prime}||_{1}$.
The cumulative regret $Reg(T)$ is defined as:
@@
where $\Delta_{S_{t}}=\max\{r(S^{*},\mu)-r(S_{t},\mu),0\}$ and $S^{*}$ is the greedy optimal action based on $\mu$, which is obtained by Algo. 3. The maximum reward gap is set as $\Delta_{\max}=\max_{S}\Delta_{S}$. The maximum reward gap of actions containing arm $s$ is $\Delta^{\max}_{s}=\max_{S:s\in S}\Delta_{S}$. We define $S^{*}$ and $S_{t}$ are sequences, i.e., $S^{*}=(s^{*}_{1},\dots,s^{*}_{K})$, $S_{t}=(s_{t,1},\dots,s_{t,K})$, and $s^{*}_{k}\in S^{*}$ is the $k$-th element added to $S^{*}$. We denote $S^{*}_{k}=(s^{*}_{1},\dots,s^{*}_{k})$ and $S_{t,k}=(s_{t,1},\dots,s_{t,k})$. For any arm $s$, define the marginal reward gap $\Delta_{s,k}=r(S^{*}_{k},\mu)-r(S^{*}_{k-1}\cup\{s\},\mu)$.
Inspired by previous work in combinatorial Thompson sampling (Chen et al., 2013; Kong et al., 2021), we provide a regret upper bound for TSAdj.
@@
(Upper Bound) Under assumptions 3.13.2 and 3.3 and with outcomes $X_{t}$ defined in Eq. 8, the regret $Reg(T)$ of TSAdj can be upper bounded by:
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}\\ &+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\\ &=O\left(\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{\Delta_{max}L^{2}\log T}{\Delta^{2}_{s,k}}\right).\end{split}$ (10)
for any $\epsilon$ such that $\forall s\not=s^{*}_{1}$ and $s\notin S^{*}_{k},\Delta_{s,k}>2BK\epsilon$, where $U$ is a universal constant.
The proof detail is provided in Sec. C.2.
@@ 4 Evaluation
Figure 3: Testing accuracy of FedRTS and different federated pruning baselines on the three CV datasets with different densities.
To validate the efficacy of FedRTS, we conduct comprehensive experiments in this section. Notably, we evaluate each experiment five times, reporting the mean results along with the standard errors. In the figures, the shaded area represents the standard errors. In the table, the best results are highlighted in purple, while the second-best results are marked in orange.
@@ 4.1 Experiment Setup
We provide a brief overview of the experimental setup in this section, with additional details in Appendix D.
Datasets and Models. Following previous work (Huang et al., 2023, 2024; Bibikar et al., 2022; Jiang et al., 2022) on federated dynamic pruning, we conduct experiments on CV tasks using two lightweight models, ResNet18 (He et al., 2016) and ShuffleNetV2 (Zhang et al., 2018), across three well-known image classification datasets: CIFAR-10 (Krizhevsky et al., 2009), CINIC-10 (Darlow et al., 2018), and SVHN (Netzer et al., 2011). For NLP tasks, we use the GPT-2-32M model on the TinyStories dataset (Eldan & Li, 2023), a language understanding benchmark designed for small language models.
Federated Learning Settings. Following the settings used in previous frameworks (Huang et al., 2023, 2024; Bibikar et al., 2022), we conduct 500 communication rounds with 5 local training epochs. Batch sizes are set to 64 for CV tasks and 16 for NLP tasks. In each round, 10 clients are randomly selected from a total of $\mathcal{N}=100$ clients. We perform an outer loop for topology adjustment every $\Delta T=10$ inner loops until the total number of loops $t$ reaches $T_{end}=300$.
Baseline Settings. We select the following SOTA federated pruning frameworks as baselines: PruneFL (Jiang et al., 2022), FedTiny (Huang et al., 2023), FedDST (Bibikar et al., 2022), FedSGC (Tian et al., 2024), and FedMef (Huang et al., 2024). These methods incorporate dynamic pruning in FL and achieve SOTA performance across various settings. We adopt all hyperparameters specified in the original papers, except for those mentioned in the previous section. Additionally, some modules are disabled due to high resource demands, such as adaptive batch normalization selection in FedTiny, as detailed in Appendix D.3.
@@
Table 1: Performance comparison on TinyStories with GPT-2-32M
Implementation Details. Following previous work (Huang et al., 2023, 2024), we exclude bias terms, normalization layers, and the output layer from pruning while ensuring the overall density $d\leq d^{\prime}$. Layer-wise sparsity follows the Erdos-Renyi-Kernel distribution (Evci et al., 2020). We set the scaling factor $\lambda=10$ and trade-off ratio $\gamma=0.5$ as default, with their impacts analyzed in Sec. E.1 and E.2. Training FLOPs and communication costs are calculated as described in Sec. F.
@@ 4.2 Performance Evaluation
To demonstrate the effectiveness of FedRTS, we conduct experiments on both computer vision (CV) and neural language (NLP) tasks under various target density levels, reporting performance and communication costs for adjustment.
Computer Vision Tasks. In the CV tasks, we conduct experiments using the ResNet18 and ShuffleNetV2 models, testing four target density levels: 50%, 30%, 20%, and 10% on CIFAR-10, CINIC-10, and SVHN datasets, reporting the accuracy and communication costs per outer loop. In some cases, FedRTS achieves similar accuracy to FedAVG at the 30% density ratio; therefore, results for the 50% density ratio are omitted. As shown in Fig. 3, FedRTS consistently outperforms all baseline methods in most scenarios in terms of both accuracy and communication efficiency. For instance, in the experiment with ResNet18 in CIFAR-10, FedRTS achieves a 5.1% improvement in accuracy over the best baseline methods while maintaining a similar communication cost (20MB). These gains are attributed to the robust adjustment, stable topology, and communication efficiency provided by TSAdj in FedRTS.
@@
Natural Language Processing Tasks. To evaluate the effectiveness of FedRTS across different domains, we conduct experiments on NLP tasks using the TinyStories dataset with the GPT-2-32M model. We test three target density levels: 50%, 30%, and 20%. Perplexity (PPL), which assesses model performance by considering the entire probability distribution in generated text, serves as the primary evaluation metric. As shown in Tab. 1, FedRTS outperforms all baselines, demonstrating its superiority in the NLP domain. Notably, FedRTS achieves better performance in perplexity than in accuracy-based evaluations, even surpassing full-size FedAVG at low-density levels. This improvement may stem from its probabilistic adjustment strategy, which enhances flexibility and adaptability during optimization. Unlike deterministic approaches, probabilistic adjustment enables FedRTS to prioritize critical parameters effectively, improving its ability to adapt to complex probability distributions. This is particularly beneficial for NLP tasks, where capturing intricate text patterns is essential, leading to lower perplexity scores and stronger performance in low-resource settings.
@@ 4.3 Robustness Analysis
To demonstrate the robustness of FedRTS in the face of heterogeneous data distributions and partial client participation, we conduct experiments with different degrees of data heterogeneity and client availability in this section.
Figure 4: Accuracy on CIFAR-10 under various degrees of data heterogeneity. The lower $a$ represents a higher non-IID degree.
Figure 5: Accuracy on CIFAR-10 with various participation ratios of clients
The Impact of Data Heterogeneity. Data heterogeneity, or non-independent and identically distributed (non-IID) data, is a critical factor affecting performance in FL. To demonstrate the robustness of FedRTS, we conduct experiments with varying degrees of data heterogeneity. By adjusting the Dirichlet distribution factor $a$ across settings $\{0.3,0.5,1.0,IID\}$, we control the degree of data heterogeneity, where lower $a$ indicates higher heterogeneity. As shown in Fig. 4, FedRTS outperforms all baselines, particularly in highly heterogeneous settings, demonstrating its effectiveness in handling non-IID data. This improvement is primarily attributed to the stable topology maintained by TSAdj in FedRTS.
The Impact of Client Availablity Client availability is another significant challenge in FL. To evaluate the robustness of FedRTS under different levels of client availability, we conduct experiments with participation ratios of 100%, 50%, 30%, 10% using a total of 10 clients. To mitigate slow convergence at lower participation levels, we adopt a high learning rate of 1 with SGD, as Adam exhibited instability with NaN values in our experiments. As shown in Fig. 5, FedRTS consistently outperforms all baselines, particularly at low participation ratios, demonstrating its adaptability and robustness in limited-client scenarios.
@@ 4.4 Ablation Study
To analyze the individual contributions of FedRTS in addressing greedy and unstable topology adjustment problems, we conduct an ablation study on the CIFAR-10 dataset with various target density levels. We consider the benefits of FedRTS to come from three aspects: probabilistic adjustment, stability, and farsighted information. To evaluate these aspects independently, we consider the following variants of FedRTS: 1) FedRTS ($\gamma=0$): This variant does not use the aggregated information to update the probability distribution. 2) FedRTS ($\gamma=1$): This variant only uses the aggregated information to update the probability distribution, which reduces stability. 3) FedRTS without TS: This variant fuses historical information via a momentum mechanism similar to FedAdam (Reddi et al., 2020) and uses deterministic adjustment to obtain the new topology instead of probabilistic adjustment. 4) FedTiny: This serves as a baseline to evaluate the overall effectiveness of FedRTS.
Figure 6: Accuracy on CIFAR-10 for FedRST and various variants.
The results in Figure 6 demonstrate the effectiveness of FedRTS. First, the performance of FedRTS and FedRTS ($\gamma=0$) is similar but better than FedRTS ($\gamma=1$), demonstrating that aggregated information suffers from instability. Including individual information can make the adjusted topology more stable and achieve better performance. Second, without deterministic adjustment, FedRTS w/o TS performs much worse than FedRTS, indicating the benefits of probabilistic adjustment. Third, FedRTS w/o TS is better than FedTiny, demonstrating the effectiveness of farsighted information. Therefore, the ablation study demonstrates the effectiveness of FedRTS in terms of probabilistic adjustment, stability, and farsighted information.
4.5 Efficiency Analysis
Training Cost. We evaluate FedRTS’s efficiency by analyzing training costs on CIFAR-10 and CINIC-10. To compare expenses, we measure total training FLOPs and communication costs required to match the testing accuracy of dense FedAVG. The accuracy is averaged over 10 rounds to mitigate variance from random factors, ensuring a more stable evaluation. As shown in Tab. 2, FedRTS achieves the testing accuracy of dense FedAVG with significantly lower costs. On CINIC-10, it achieves 48% training FLOPs and the communication cost compared to the best baseline FedMef, which highlights the efficiency of FedRTS.
CIFAR-10 CINIC-10
FLOPs Comm. Cost FLOPs Comm. Cost
FedDST 3.4E16 221.9 GB 8.2E16 290.7 GB
FedMef 3.5E16 227.6 GB 6.1E16 222.3 GB
FedRTS 3.1E16 204.2 GB 2.9E16 105.6 GB
Table 2: The overall training FLOPs and communicational cost (Comm. Cost) for FedRTS and other frameworks to reach the testing accuracy of the dense FedAVG.
Convergence Behavior. Convergence is crucial for effective learning, precise predictions, and optimized resource use. To evaluate this, we track the validation accuracy of FedRTS and SOTA frameworks on CIFAR-10 against cumulative communication and computational costs, as shown in Fig. 7 and Sec. E.4. Accuracy is averaged over 10 rounds to reduce variance and provide a stable representation. Results show that FedRTS converges faster than other methods, achieving higher final accuracy with improved communication efficiency, highlighting the benefits of TSAdj in FedRTS in enhancing convergence.

### v3

@@ 3.2 Proposed FedRTS framework
3.3 Overhead Analysis
The overhead of TSAdj arises from maintaining and sampling the Beta distribution on the server. In each round, the server performs the following operations with $\langle C_{t}\rangle$ sampled clients: weight averaging with a time complexity of $\mathcal{O}(\langle C_{t}\rangle\langle W\rangle)$, computing the outcome using Quicksort with an expected time complexity of $\mathcal{O}(\langle C_{t}\rangle\langle W\rangle)$, element-wise Beta distribution updates in $\mathcal{O}(\langle W\rangle)$, and arms selection in $\mathcal{O}(\langle W\rangle)$. The overall complexity thus simplifies to $\mathcal{O}(\langle C_{t}\rangle\langle W\rangle)$. Compared to the baseline complexities of FedMef, FedTiny, and PruneFL, which are also $\mathcal{O}(\langle C_{t}\rangle\langle W\rangle)$, TSAdj achieves the same computational order. This demonstrates that TSAdj introduces no significant additional computational burden on the server, even with large models.
Critically, TSAdj introduces no additional computations on client devices compared to the federated pruning baseline, as its overhead is entirely server-side. In cross-device federated learning, the server is typically assumed to possess substantial computational resources; therefore, TSAdj only introduces a slight overhead on the system.
@@ 3.4 Theoretical Analysis
For simplifications, we omit the inner loop of FedRTS to isolate the TSAdj mechanism and consider only semi-outcomes, i.e., ignore the outcomes $X_{t,i}$ that $i\notin S_{t}$. We adopt the following assumptions:
@@
[…]
@@
(Independent importance score) The importance scores used for pruning are treated as mutually independent.
Assumption 3.3.
(Mean-field Approximation) when $\{x_{i}\}$ are mutually independent, the discriminative function $h(\cdot)$ admits a mean-field approximation: $h(i,x,\kappa)=\mathbf{1}_{i\in\{x,\kappa\}}\approx\mathbf{1}_{x_{i}\geq\sigma}$, where the threshold $\sigma$ is determined self-consistently via mean-field decoupling (independent of individual $x_{i}$) and satisfies $\langle x\rangle-\sum_{i}^{\langle x\rangle}\mathrm{CDF}_{i}(\sigma)=\kappa$, where $\mathrm{CDF}_{i}$ is the cumulative distribution function of $x_{i}$.
Assumption 3.1 is standard in bandit optimization work [26]. Assumption 3.2 is implicitly or explicitly adopted in pruning methods [46, 12, 52]. Assumption 3.3 holds for large scale $\langle x\rangle$ (>100k) [53, 54, 35].
While weight magnitudes are not strictly theoretically independent when used as importance scores, both empirical and theoretical evidence supports simplification in Assumption 3.2: (1) Recent work demonstrates that trained weights in large networks exhibit asymptotic independence [54, 53], and (2) this assumption is well-established in the pruning work [12, 46], where weight dependencies are typically disregarded during selection. While other importance scores (e.g., Fisher information [52]) might better satisfy independence, they are computationally prohibitive in FL. Thus, we follow prior works [24, 19] to apply weight magnitude pruning, acknowledging a minor discrepancy between theory and practice. Under assumptions 3.2 and 3.3, outcomes $X_{t,i}$ simplify to:
@@
which is mutually independent.
To analyze the worst-case performance of TSAdj, we define the cumulative regret $Reg(T)$ as:
@@
where $\Delta_{S_{t}}=\max\{r(S^{*},\mu)-r(S_{t},\mu),0\}$ and $S^{*}$ is the greedy optimal action based on $\mu$, which is obtained by Algorithm 3. The maximum reward gap is set as $\Delta_{\max}=\max_{S}\Delta_{S}$. The maximum reward gap of actions containing arm $s$ is $\Delta^{\max}_{s}=\max_{S:s\in S}\Delta_{S}$. We define $S^{*}$ and $S_{t}$ are sequences, i.e., $S^{*}=(s^{*}_{1},\dots,s^{*}_{K})$, $S_{t}=(s_{t,1},\dots,s_{t,K})$, and $s^{*}_{k}\in S^{*}$ is the $k$-th element added to $S^{*}$. We denote $S^{*}_{k}=(s^{*}_{1},\dots,s^{*}_{k})$ and $S_{t,k}=(s_{t,1},\dots,s_{t,k})$. For any arm $s$, define the marginal reward gap $\Delta_{s,k}=r(S^{*}_{k},\mu)-r(S^{*}_{k-1}\cup\{s\},\mu)$. Inspired by previous work [5, 26], we provide a regret upper bound for TSAdj.
@@
(Upper Bound) Under assumptions 3.23.3 and 3.1 and with outcomes $X_{t}$ defined in Eq. 7, the regret $Reg(T)$ of TSAdj can be upper bounded by:
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\\ &=O\left(\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{\Delta_{max}L^{2}\log T}{\Delta^{2}_{s,k}}+\Delta_{max}\right).\end{split}$ (9)
for any $\epsilon$ such that $\forall s\not=s^{*}_{1}$ and $s\notin S^{*}_{k},\Delta_{s,k}>2LK\epsilon$, where $U$ is a universal constant.
The proof is provided in Sec. C.2. Crucially, our theorem hold for any pruning method satisfying Assumption 3.2 and 3.3, including future advancements in efficient independent-scoring techniques.
@@ 4 Evaluation
Figure 3: Testing accuracy of FedRTS and different federated pruning baselines on the four CV datasets with different densities, where the ratio is communication cost relative to dense FedAVG.
To validate the efficacy of FedRTS, we conduct comprehensive experiments in this section. Notably, we evaluate each experiment five times, reporting the mean results along with the standard errors. In the figures, the shaded area represents the standard errors. In the table, the best results are highlighted in purple color, while the second-best results are marked in blue color.
@@ 4.1 Experiment Setup
We provide an overview of the experimental setup in this section, with additional details in Sec. G.
Datasets and Models. Following previous work [18, 19, 1, 24] on federated dynamic pruning, we conduct experiments on CV tasks using two lightweight models, ResNet18 [13] and ShuffleNetV2 [66], across four well-known image classification datasets: CIFAR-10 [27], CINIC-10 [7], TinyImageNet [8] and SVHN [43]. For NLP tasks, we use the GPT-2-32M model on the TinyStories dataset [10], a language understanding benchmark designed for small language models.
Federated Learning Settings. Following the settings used in previous frameworks [18, 19, 1], we conduct $T_{max}=500$ communication rounds with 5 local training epochs. Batch sizes are set to 64 for CV tasks and 16 for NLP tasks. In each round, 10 clients are randomly selected from a total of $\mathcal{N}=100$ clients. To generate data heterogeneity on the client, we utilize the Dirichlet distribution with factor $a$. The types of heterogeneity for CV and NLP tasks are different. CV tasks apply label imbalance with $a=0.5$ by default. Unlabeled NLP tasks set data quantity imbalance, with $a=5$ by default. We perform an outer loop for topology adjustment every $\Delta T=10$ inner loops until the total number of loops $t$ reaches $T_{end}=300$. The optimizer used is SGD, with a learning rate of $\eta=0.01$.
Baseline Settings. We select the following SOTA federated pruning frameworks as baselines: PruneFL [24], FedTiny [18], FedDST [1], FedSGC [57], and FedMef [19]. These methods incorporate dynamic pruning in FL and achieve SOTA performance across various settings, as detailed in Sec. G.2.
Implementation Details. Following previous work [18, 19], we exclude bias terms, normalization layers, and the output layer from pruning while ensuring the overall density $d\leq d^{\prime}$. Layer-wise sparsity follows the Erdos-Renyi-Kernel distribution [11]. We set the scaling factor $\lambda=10$ to match the number of selected clients per loop, ensuring that the averaged information maintains individual-level influence. The trade-off ratio is set as $\gamma=0.5$ to assign equal importance to the individual and aggregated model. We set the layer-wise $\kappa^{l}=K^{l}-\frac{\alpha_{adj}}{2}(1+\cos{(\frac{t\pi}{T_{end}})})K^{l}$, where $K^{l}$ is the number of active weights at $l$-th layer and $\alpha_{adj}$ is 0.4 by default, is adopted directly from prior works [19, 11]. Training FLOPs and communication costs are calculated as described in Sec. G.4.
@@
Table 1: Performance comparison on TinyStories with GPT-2-32M.
Figure 4: Accuracy on CIFAR-10 under various degrees of data heterogeneity.
@@ 4.2 Performance Evaluation
To show the effectiveness of FedRTS, we conduct experiments on CV and NLP tasks under various target density levels. Broader experiments including efficiency analysis are in Sec. F.2 and F.3.
Computer Vision Tasks. In the CV tasks, we conduct experiments using the ResNet18 and ShuffleNetV2 models, testing four target density levels: 50%, 30%, 20%, and 10% on CIFAR-10, CINIC-10, TinyImageNet, and SVHN datasets, reporting the accuracy and communication costs ratio per outer loop. In some cases, FedRTS achieves similar accuracy to FedAVG at the 30% density ratio; therefore, results for the 50% density ratio are omitted. As shown in Fig. 3, FedRTS consistently outperforms all baseline methods in most scenarios in terms of both accuracy and communication efficiency. For instance, in the experiment with ResNet18 in CIFAR-10, FedRTS achieves a 5.1% improvement in accuracy over the best baseline methods while maintaining a similar communication cost. These gains are attributed to the robust adjustment, stable topology, and communication efficiency provided by TSAdj in FedRTS.
@@
Natural Language Processing Tasks. To evaluate the effectiveness of FedRTS across different domains, we conduct experiments on NLP tasks using the TinyStories dataset with the GPT-2-32M model. We test three target density levels: 50%, 30%, and 20%. Perplexity (PPL), which assesses model performance by considering the entire probability distribution in generated text, serves as the primary evaluation metric. As shown in Tab. 1, FedRTS outperforms all baselines, demonstrating its superiority in the NLP domain. Notably, FedRTS achieves better performance in perplexity than in accuracy-based evaluations, even surpassing full-size FedAVG at low-density levels. This improvement may stem from its probabilistic adjustment strategy, which enhances flexibility and adaptability during optimization. Unlike deterministic approaches, probabilistic adjustment enables FedRTS to prioritize critical parameters effectively, improving its ability to adapt to complex probability distributions. This is particularly beneficial for NLP tasks, where capturing intricate text patterns is essential, leading to lower perplexity and stronger performance in low-resource settings.
@@ 4.3 Robustness Analysis
To demonstrate the robustness of FedRTS in the face of heterogeneous data distributions and partial client participation, we conduct experiments with different degrees of data heterogeneity and client availability in this section. Further experiments on the impact of the hyperparameters $\gamma$, $\lambda$, $\kappa$, and $\Delta T$ are provided in Sec. F.1.
Figure 5: Accuracy on CIFAR-10 with various participation ratios of clients.
Figure 6: Accuracy on CIFAR-10 for FedRTS and various variants.
The Impact of Data Heterogeneity. Data heterogeneity, or non-independent and identically distributed (non-IID) data, is a critical factor affecting performance in FL. To demonstrate the robustness of FedRTS, we conduct experiments with varying degrees of data heterogeneity on the CIFAR-10 and TinyStories datasets. By adjusting the Dirichlet distribution factor $a$, we control the degree of data heterogeneity, where a lower $a$ indicates higher heterogeneity. As shown in Fig. 4 and Fig. 14, FedRTS outperforms all baselines, particularly in highly heterogeneous settings, demonstrating its effectiveness in handling non-IID data. This improvement is primarily attributed to the stable topology produced by TSAdj in FedRTS.
The Impact of Client Availability. Client availability is another significant challenge in FL. To evaluate the robustness of FedRTS under different levels of client availability, we conduct experiments with participation ratios of 100%, 50%, 30%, 10% using a total of 10 clients. To mitigate slow convergence at lower participation levels, we adopt a high learning rate of 1 with SGD, as Adam exhibited instability with NaN values in our experiments. As shown in Fig. 6, FedRTS consistently outperforms all baselines, particularly at low participation ratios, demonstrating its adaptability and robustness in limited-client scenarios.
@@ 4.4 Ablation Study
To analyze the individual contributions of FedRTS in addressing greedy and unstable topology adjustment problems, we conduct an ablation study on the CIFAR-10 dataset with various target density levels. We consider the benefits of FedRTS to come from three aspects: probabilistic decision, farsighted information, and stability from the fusion mechanism in Eq. 3. To evaluate these aspects independently, we consider the following variants of FedRTS: (1) FedRTS ($\gamma=0$): This variant does not use the aggregated information to update the probability distribution. (2) FedRTS ($\gamma=1$): This variant only uses the aggregated information to update the probability distribution, which disables the fusion mechanism to reduce stability. (3) FedRTS without TS ($\gamma=1$): This variant fuses historical information via a momentum mechanism similar to FedAdam [49] and uses deterministic adjustment to obtain the new topology, disabling both probabilistic decision and fusion mechanism. (4) FedTiny: This serves as a baseline to evaluate the overall effectiveness of FedRTS, disabling all three aspects.
The results in Fig. 6 demonstrate the effectiveness of FedRTS. First, the performance of FedRTS and FedRTS ($\gamma=0$) is similar but better than FedRTS ($\gamma=1$), demonstrating that aggregated information suffers from instability. A fusion mechanism including individual information can make the adjusted topology more stable and achieve better performance. Second, without deterministic adjustment, FedRTS w/o TS ($\gamma=1$) performs much worse than FedRTS ($\gamma=1$), indicating the benefits of probabilistic adjustment. Third, FedRTS w/o TS ($\gamma=1$) is better than FedTiny, showing the effectiveness of farsighted information. Thus, the ablation study demonstrates FedRTS’s effectiveness in terms of probabilistic decision, farsighted information, and stability from the fusion mechanism.